OpenStudy (anonymous):

is this correct? problem => (1/8x^2 - 1/4x - 3/8) + ( -1/2x^2 - 1/4x - 1/2) my answer is => 17.5x^2 - 1/2x - 7/8

OpenStudy (anonymous):

Check your first term

OpenStudy (anonymous):

should it be -17.5 is that what ur asking?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Everything other than the first term is correct. 1/8 - 4/8 = -3/8

OpenStudy (anonymous):

\[\frac{1}{8}-\frac{1}{2}\] Common denominator is 8 \[\frac{1}{8}-\frac{1*4}{2*4}\]

OpenStudy (anonymous):

i am lost then what are the steps? tonight, has not been a math noght (smiling)

OpenStudy (anonymous):

and what @Calcmathlete got... :)

OpenStudy (anonymous):

lol we just posted it... everything else is corrected just ignore the variable for a second and focus on the fractions... and work them like your working regular numbers and then plug back in your variables

OpenStudy (anonymous):

\[(\frac{1}{8}x^2 - \frac{1}{4}x - \frac{3}{8}) + ( -\frac{1}{2}x^2 - \frac{1}{4}x - \frac{1}{2}) \] Okay so.. -1/4-1/4 = -2/4 and -3/8-1/2 =\[-\frac{3}{8}-\frac{1*4}{2*4}\] =\[-\frac{3}{8}-\frac{4}{8}\] =\[-\frac{7}{8}\]

OpenStudy (anonymous):

boy am i way off..

OpenStudy (anonymous):

Do you understand where you went wrong @fuzzyspawn though?

OpenStudy (anonymous):

y did the calculator give me the numbers i put down?

OpenStudy (anonymous):

i wont be using the calculator anymore i will go back to long hand.

OpenStudy (anonymous):

I was using the polinomial calculator from online someone reccommended

OpenStudy (anonymous):

polynomial

OpenStudy (anonymous):

Hmm, I don't trust my calculator sometimes with fractions... did you use a lot of brackets because some calculators require more than others. Try putting it in like this and let me know what you get (1/8)-(1/2)

OpenStudy (anonymous):

No... it might cost you to loose time on exams, test and quizzes.

OpenStudy (anonymous):

Oh... hmm, well computer's do have some faults so Idk

OpenStudy (anonymous):

\[(1/8x{}^{\wedge}2-1/4x-3/8)+(-1/2x{}^{\wedge}2-1/4x-1/2)=-\frac{3 x^2}{8}-\frac{x}{2}-\frac{7}{8} \]

OpenStudy (anonymous):

But that should only be used as a check have you tried wolfram alpha? I in most instances don't go wrong there.

OpenStudy (anonymous):

i havent heard about that one site, is it good?

OpenStudy (anonymous):

Yes it is...

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (anonymous):

cool then i will try it I appreciate your help! getting (math headache) gonna get back later thanks

OpenStudy (anonymous):

Okay, no problemo!